### v2

# Odo[Uncaptioned image]: Depth-Guided Diffusion for Identity-Preserving Body Reshaping

Abstract
Human shape editing enables controllable transformation of a person’s body shape, such as thin, muscular, or overweight, while preserving pose, identity, clothing, and background. Unlike human pose editing, which has advanced rapidly, shape editing remains relatively underexplored. Current approaches typically rely on 3D morphable models or image warping, often introducing unrealistic body proportions, texture distortions, and background inconsistencies due to alignment errors and deformations. A key limitation is the lack of large-scale, publicly available datasets for training and evaluating body shape manipulation methods. In this work, we introduce the first large-scale dataset of 18,573 images across 1523 subjects, specifically designed for controlled human shape editing. It features diverse variations in body shape, including fat, muscular and thin, captured under consistent identity, clothing, and background conditions. Using this dataset, we propose Odo, an end-to-end diffusion-based method that enables realistic and intutive body reshaping guided by simple semantic attributes. Our approach combines a frozen UNet that preserves fine-grained appearance and background details from the input image with a ControlNet that guides shape transformation using target SMPL depth maps. Extensive experiments demonstrate that our method outperforms prior approaches, achieving per-vertex reconstruction errors as low as 7.5mm, significantly lower than the 13.6mm observed in baseline methods, while producing realistic results that accurately match the desired target shapes.
Figure 1: Odo enables realistic body shape transformation by adjusting semantic sliders (e.g., weight) that directly modify the SMPL model. The modified SMPL then guides the transformation of the input image. Natural language prompts like “make the person muscular,” “fatter,” or “thinner” further refine the output. In addition to the weight parameter, in the last column we make the person muscular by giving the prompt “Make the person muscular”. The model supports accurate shape edits from a single reference image.
1 Introduction
Digital human shape editing is a critical task, enabling realistic manipulation of body shapes across various domains, such as entertainment, virtual fashion, health visualization, and personalized digital avatars. Moreover, it can support medical interventions by visualizing healthy body shapes for conditions such as anorexia and bulimia nervosa [31, 7]. Human shape editing involves modifying a person’s body size, shape, and proportions in images, making them appear thinner, heavier, or more muscular while preserving key attributes like pose, identity, clothing, and background.
Previous methods for human shape editing have primarily relied on 3D morphable models or image warping techniques. However, these methods often introduce visual artifacts and background distortions, especially under large deformations, due to their reliance on explicit geometric correspondences. In contrast, human pose editing has seen substantial progress, largely due to the availability of large-scale pose editing datasets [20]. Capturing diverse poses is straightforward, as an individual can rapidly shift poses during short capture sessions, or poses can be automatically extracted from video footage. However, gathering data of the same person exhibiting multiple distinct body shapes presents significant practical challenges. Thus, the absence of large-scale datasets with identity-consistent body shape variations remains a critical barrier to developing robust shape editing techniques.
To address this, we introduce ChangeLing18K, the first large-scale dataset specifically curated for human shape editing. It features diverse body types including thin, muscular, overweight, with each pair depicting the same subject across different shapes while maintaining consistent identity, pose, clothing, and viewpoint. This facilitates precise and identity-preserving model training. Building on this dataset, we propose Odo, to the best of our knowledge, the first ever end-to-end diffusion-based framework designed specifically for human shape editing. Recently, diffusion models have demonstrated remarkable success in different image-editing tasks [23, 6, 22, 16]. Their iterative denoising mechanism, combined with strong priors learned from large datasets, enables fine-grained control and high-quality outputs. However, traditional methods like text-based or point-dragging-based methods [3, 39] and are not capable of human body shape transformation. These techniques lack the precise control for editing human body shapes often causing artifacts or undesired results.
Odo achieves precise, linear, and scalable shape control through the integration of explicit shape guidance into the diffusion process through a combination of a frozen SDXL UNet [30] and a depth ControlNet-conditioned [45] diffusion model. The frozen UNet preserves identity, appearance, and background details from the input image, while SMPL-based depth maps condition the diffusion model towards achieving accurate and visually coherent target body shapes. Users can interactively adjust semantic body attributes through intuitive sliders, facilitating realistic and precise reshaping without compromising visual fidelity.
We also introduce the first comprehensive benchmark specifically designed to quantitatively evaluate human shape editing performance. Our benchmark provides metrics assessing both image quality relative to ground truth and the accuracy of target body shape transformations. Experimental results demonstrate that Odo outperforms existing methods across all evaluated metrics by a large margin.
In summary, our contributions are threefold: (1) ChangeLing18K, the first large-scale dataset specifically designed for human shape editing; (2) Odo, a novel diffusion-based model capable of realistic and precise shape transformations trained on our dataset; and (3) a comprehensive benchmark for quantitatively evaluating human shape editing. We will publicly release our dataset and model to promote future research and reproducibility in human body reshaping.
Figure 2: Interactive UI with semantic sliders for shape editing. In this example, the semantic sliders are adjusted to modify the body shape, resulting in the person appearing significantly fatter compared to the original input image.
2 Related Work
Human body shape editing requires specialized approaches beyond general image editing, as it demands precise anatomical transformations of the human body. In this section we discuss the evolution of human-shape editing methods, highlight general diffusion-based techniques, and the current scope of SMPL conditioning in diffusion models.
Human Shape Editing. Human body shape editing has evolved through several approaches, each with distinct advantages and limitations.
IDENTITY GENERATIONRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionAI Generated+Real FacesIDENTITY CONDITIONEDFULL-BODY GENERATIONPuLIDTransform-basedFull-bodyDescriptionsRefer to captionRefer to captionTHINFATBACKGROUND & MODELASPECT-RATIO NORMALISATIONFLUX.1 KONTEXT-PRORefer to caption(BACKGROUND) RT-DETRv2 SAM 2.1 +Refer to captionRefer to caption(MASKS)SCALING&PASTINGExtractReference MaskSize & PositionRefer to captionTHINRefer to captionFAT
Figure 3: End-to-end dataset generation pipeline. Identities are created using FLUX.1-dev[18] and license-free face images. PuLID[9], a diffusion-based tuning-free ID customization method for text-to-image generation, generates thin, fat, and muscular variations for each identity. Human masks from RT-DETRv2[24]+SAM2.1[32] are scaled and pasted onto a common FLUX-Kontext[19] refined background for consistency.
Zhou et al. [47] pioneered parametric body reshaping by fitting 3D morphable body model (SCAPE [1]) via user-guided pose alignment and contour-based shape fitting. Semantic edits applied to the fitted model are transferred back to the image through body-aware warping. MovieReshape [13] by Jain et al. extended this by automating the fitting process, enabling temporally coherent reshaping in videos. Richter et al. [34] introduced a real-time reshaping system using Kinect sensors, employing automatic shape initialization, pose tracking, segmentation refinement, and foreground-specific image warping. Despite their effectiveness, these 3D model-based approaches require manual intervention and generalize poorly beyond constrained datasets. They also suffer from fitting inaccuracies propagating into noticeable warping artifacts, pose tracking instabilities under occlusions, and also unintended global warping effects on closely positioned body parts.
Ren et al. [33] introduced a method that predicts pixel-level deformation flows guided by body structure using skeletons and Part Affinity Fields. Their Structure Affinity Self-Attention ensures consistent changes across related body parts by leveraging visual and structural relationships. However, this approach struggles with significant shape edits, causing unnatural warping and background distortions.
Okuyama et al. [28] combine SMPL-X parametric body model initialization with diffusion-based refinement for pose and shape editing. They fit and modify the 3D model, followed by a two-stage diffusion refinement with weak noise to preserve identity and remove projection artifacts. However, their approach requires per-subject fine-tuning and careful noise schedule tuning. It also struggles with loose clothing and background distortions due to projection of the original image onto the edited SMPL model. Consequently, these methods remain ineffective for robust human body reshaping, often resulting in unrealistic outputs with background distortions.
Image Editing using Diffusion. Foundational work such as Denoising Diffusion Probabilistic Models (DDPM) [11] and Latent Diffusion Models (LDM) [35] introduced scalable frameworks that enable controllable and high-resolution image synthesis. These models have served as the foundation for numerous diffusion-based image-editing works across a broad range of tasks, including inpainting, colorization, semantic modification, and subject personalization.
Methods such as Palette [36] and SDEdit [25] have demonstrated the capability of diffusion models for image-to-image tasks like colorization and sketch-to-image synthesis. Prompt-to-Prompt [10] provided fine-grained attention control, enabling localized edits without retraining. Building upon these techniques like, Imagic [15] and InstructPix2Pix [3] introduced semantically rich, text-driven image editing. Null-text inversion [26] proposed an efficient inversion scheme for real images by optimizing the unconditional textual embedding used in classifier-free guidance, enabling accurate reconstruction and intuitive prompt-driven editing without fine-tuning the model weights. Meanwhile, works like DragDiffusion [39] and DragonDiffusion [27] enabled point-based image editing by allowing users to manipulate images by dragging handle points to target points.
These methods are designed for semantic, structural, or stylistic editing but are unsuitable for large-scale anatomical transformations in the human body while preserving fine-grained identity and clothing details. Human shape editing requires manipulating body proportions in a controllable manner, beyond the capacity of latent inversion or prompt-conditioned editing.
Table 1: Comparison between BR-5K and Our Body Shape Transformation Dataset. Qualitative comparison is provided in the supplementary material.
Feature BR-5K Our Dataset
# of Pairs 5,000 18.573
Body Shape Variation Extent Minimal Wide
Identity Consistency - Yes
% of Female Pairs 92.4 44.2
% of Male Pairs 7.6 55.8
SMPL Conditioned Diffusion. Recent methods integrate SMPL models into diffusion frameworks to enhance spatial and anatomical consistency in human image generation. CHAMP [48] utilizes SMPL-derived spatial representations as conditioning inputs to a latent diffusion model. By aligning these SMPL conditions with a reference image’s shape parameters, CHAMP achieves temporally coherent human animations. However, its scope is limited to pose variation and does not support body shape editing. Buchheim et al. [4] propose an SMPL-conditioned ControlNet for text-to-image generation, explicitly encoding SMPL shape and pose parameters into the diffusion UNet cross-attention. However, their method requires synthetic-to-real domain adaptation and does not directly facilitate shape editing. These methods underscore the potential of SMPL conditioning to enhance human shape control within diffusion models.
3 ChangeLing18K Dataset
Despite substantial progress in image generation, large-scale, high-resolution datasets tailored specifically for human body shape editing remain scarce. Existing datasets typically lack controlled shape variations of the same individual under consistent identity, pose, clothing, and background conditions. For instance, the BR-5K dataset [33] provides 5,000 high-quality portrait pairs with professionally retouched versions; however, each image has the face-blurred out and includes only a single minimal transformation, which is often barely noticeable (see Table 1). To address this gap, we introduce ChangeLing18K, a novel large-scale dataset specifically designed for identity-consistent human shape transformations.
3.1 Dataset Generation Workflow
ChangeLing18K is synthesized through a four–stage, model–driven pipeline that guarantees identity fidelity, pose and attire consistency, and high–resolution realism across three body–types (thin, fat, muscular) Fig. 4. The complete dataset generation pipeline is shown in Fig. 3 and detailed below.
1. Identity Creation: We initialize each synthetic subject with a high–resolution facial portrait generated by FLUX.1-dev, a 12-B parameter rectified–flow diffusion transformer that excels at photorealistic face synthesis from textual prompts [18]. To increase demographic diversity, we additionally sample license–free reference faces from Pexels [29] and Unsplash [40]. In total, we create a corpus of 1523 human face images.
[…]
Thin Fat Muscular
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 4: Example dataset images showcasing subjects in thin, fat, and muscular forms. The dataset includes such images with diverse backgrounds, poses, clothing, and body transformations.
Figure 5: The left figure illustrates our training pipeline, where the SMPL depth-map extracted from the target image conditions the ReshapeNet via ControlNet for the desired shape transformation. Features from the reference image, extracted by the ReferenceNet, are integrated into the ReshapeNet using spatial self-attention. The right figure depicts the inference pipeline, where SMPL parameters are initialized from the input image, adjusted based on semantic attributes, and rendered into a target SMPL depth map. This depth map, along with input image features, conditions the ReshapeNet during inference to generate the final transformed output through iterative denoising.
3. Background and Aspect–ratio Normalization: The three body type images for each identity sometimes exhibit slight shifts in background alignment and minor differences in perceived subject size. This variation arises due to the inherent stochastic nature of diffusion models: each generation begins from a new Gaussian noise seed, and even subtle alterations in the input prompt—such as changing only the body type description—can influence the model’s latent-space trajectories, resulting in these visual discrepancies. To eliminate this variability we carry out a lightweight, fully automatic post-process:
i) Reference Image selection. The thin image is treated as the baseline reference because it’s smaller silhoutte exposes the most background, yielding a neear complete scene for subsequent background completion step.
ii) Subject Segmentation. We first detect the person in the reference image using RT-DETRv2[24], then pass the resulting bounding box to SAM 2.1[32] to produce a binary mask. The same detector–segmenter pipeline is used to obtain masks for the remaining two body types.
iii) Background Completion. We pass the thin reference image to FLUX.1 Kontext [pro] [19] prompting it to remove the person and complete the background. This yields a clean background image shared by the entire triplet (thin, fat, muscle).
iv) Uniform Scaling & Composition. From the thin reference mask we measure the height and record the bottommost coordinates of the mask. Each non-thin mask is resized uniformly to match it’s height, maintaining aspect ratio, and composited onto the clean background at the same coordinates. This ensures consistent framing across images while preserving real shape variations.
The resulting transformation triplets (thin, fat and muscle) have an identical background and uniform subject scale. This removes irrelevant variations that could negatively affect subsequent training or evaluation.
[…]
While minor pose variations exist across shape transformations for each identity, our model remains robust to these differences.
4 Method
Given the above dataset we propose Odo, a diffusion-based approach for reshaping humans in images using target SMPL depth maps. The SMPL model [21] represents the human body’s 3D shape with parameters $\boldsymbol{\beta}\in\mathbb{R}^{10}$ and pose with joint rotation parameters $\boldsymbol{\theta}\in\mathbb{R}^{72}$. To generate the conditioning depth maps, we first fit the SMPL model to each individual in our dataset using the method introduced in [37]. The optimized $\boldsymbol{\beta}$ and $\boldsymbol{\theta}$ parameters are subsequently used to render spatially aligned depth maps.
During training, we use paired images from our dataset of the same individual, captured with consistent clothing and background, but exhibiting different body shapes. Given an input image $\boldsymbol{I}$ and an SMPL target depth map $\boldsymbol{S}$, our objective is to transform the person in $\boldsymbol{I}$ to match the shape in $\boldsymbol{S}$.
For inference on in-the-wild images without ground-truth SMPL depth maps, we use a mapping between semantic attributes and SMPL $\boldsymbol{\beta}$ shape parameters [8]. This mapping enables modification of semantic attributes like weight, height, and body size to control the SMPL $\boldsymbol{\beta}$ shape parameters. Adjusting these parameters modifies the depth map, ensuring the generated output image aligns with the desired body shape. We show this interface in Fig. 2.
4.1 Model Architecture
Our model (Fig. 5) consists of a ReshapeNet supported by three key modules: a ReferenceNet, an IP-Adapter [44], and a Depth ControlNet [45]. The ReferenceNet captures detailed features (background, clothing, identity) from input image $\boldsymbol{I}$ and feeds them to the ReshapeNet. The IP-Adapter provides high-level feature assistance, while the Depth ControlNet performs SMPL conditioning for body shape transformation. We discuss each module and their role in more detail below.
ReferenceNet. To retain fine details of the input image, the IP-Adapter alone is insufficient. Following previous works [12, 43, 5, 48], we incorporate a frozen SDXL UNet to extract intermediate image features. These intermediate latents are injected into the self-attention blocks of the ReshapeNet by concatenating them with the hidden states before performing spatial self-attention. This operation is formulated as:
$\text{Attn}(\mathbf{Q^{\prime}},\mathbf{K^{\prime}},\mathbf{V^{\prime}})=\text{Softmax}\left(\frac{\mathbf{Q^{\prime}}\mathbf{K^{\prime}}^{T}}{\sqrt{d}}\right)\mathbf{V^{\prime}}$ (1)
where $\mathbf{Q^{\prime}}=\mathbf{W_{q}}[z,y]$, $\mathbf{K^{\prime}}=\mathbf{W_{k}}[z,y]$, and $\mathbf{V^{\prime}}=\mathbf{W_{v}}[z,y]$. Here, $z$ denotes the intermediate latents of the ReshapeNet, $y$ represents the corresponding intermediate latents from the frozen UNet encoder, and $[z,y]$ indicates concatenation. We pass only the first half of the attention output to subsequent layers. The ReferenceNet uses a fixed prompt “A photo of a person” throughout training.
Image Prompt Adapter. To condition the diffusion model with high-level features from the input image, we employ the IP-Adapter module [44]. The IP-Adapter encodes the input image using a CLIP image encoder, then projects the features through linear and self-attention layers to generate image embeddings. These embeddings are integrated into the ReshapeNet via cross-attention, analogous to text embeddings. The combined cross-attention operation is:
$Z^{\text{new}}=\text{Attn}(\mathbf{Q},\mathbf{K_{t}},\mathbf{V_{t}})+\text{Attn}(\mathbf{Q},\mathbf{K_{i}},\mathbf{V_{i}})$ (2)
where $\mathbf{Q}$ is the query matrix from the ReshapeNet’s intermediate representation, $\mathbf{K_{t}}$ and $\mathbf{V_{t}}$ are the key and value matrices from text embeddings, and $\mathbf{K_{i}}$ and $\mathbf{V_{i}}$ are the key and value matrices from image embeddings.
SMPL Depth ControlNet. ControlNets [45] enable conditioning of text-to-image diffusion models using spatial inputs such as Canny edges, normal maps, and depth maps. In Odo, we integrate a Depth ControlNet that applies conditioning to the middle and decoder blocks of the ReshapeNet through residual connections. The conditioning depth map is rendered from the target SMPL parameters and is aligned with the target image.
ReshapeNet. The ReshapeNet serves as the base UNet in our model, built upon the SDXL UNet architecture [30]. During training, the target image is encoded through a variational autoencoder to obtain latents $\boldsymbol{z_{0}}$. Noise is progressively added to produce $\boldsymbol{z_{t}}$ at timestep $t$, which serves as input to the ReshapeNet along with features from the ControlNet and ReferenceNet. The UNet output is then decoded to produce the denoised image.
To guide category-specific transformations, we incorporate targeted text prompts based on the transformation type: “Make the person fatter”, “Make the person thinner”, or “Make the person muscular”. These prompts provide crucial semantic control, particularly for muscularity transformations where SMPL-derived depth maps are insufficient. While depth maps effectively capture coarse shape attributes like overall body size and proportions, they cannot encode fine-grained features such as muscle definition. Prompt-based conditioning compensates for this limitation by introducing high-level semantic intent, enhancing the model’s ability to generate accurate and detailed body modifications. We demonstrate the effectiveness of these prompts in Sec. 5.3.
4.2 Training Setup
Given an input reference image $\boldsymbol{I}$, target SMPL depth map $\boldsymbol{S}$, and text prompt $\boldsymbol{c_{t}}$, the model learns to predict noise added to the noised latent $\boldsymbol{z_{t}}$ using the standard diffusion learning objective:
$\mathcal{L}=\mathbb{E}_{\boldsymbol{z_{0}},t,\boldsymbol{c_{t}},\boldsymbol{I},\boldsymbol{S},\epsilon\sim\mathcal{N}(0,1)}\left[\|\epsilon-\epsilon_{\theta}(\boldsymbol{z_{t}},t,\boldsymbol{c_{t}},\boldsymbol{I},\boldsymbol{S})\|_{2}^{2}\right]$ (3)
where $\epsilon_{\theta}$ is the model’s noise prediction parameterized by trainable parameters $\theta$.
5 Experiments
We conduct comprehensive experiments to evaluate our approach, including qualitative and quantitative comparisons with existing methods across diverse test cases, along with ablation studies to validate key design choices.
5.1 Implementation Details
We train Odo on our training set combined with a subset of the DeepFashion-MultiModal dataset [14], totaling 20,000 image pairs. The DeepFashion-MultiModal images provide exposure to complex garments and varied faces, with images paired with themselves during training. All SMPL depth maps are precomputed to minimize memory usage and enhance training efficiency.
[…]
For initialization, the ReshapeNet uses pre-trained SDXL UNet weights and is fine-tuned jointly with the IP-Adapter (initialized from its pre-trained checkpoint). The ReferenceNet is initialized with SDXL weights and kept frozen during training. The ControlNet uses pre-trained depth ControlNet weights [42] and also remains frozen throughout training.
Our model requires approximately 23GB GPU memory and 18 seconds for single-image inference.
Figure 6: Odo generates realistic body transformations while preserving identity, clothing, and background details. Examples show transformations from original images to (a) thinner, (b) overweight, and (c) muscular body types across various poses including sitting and standing positions.
5.2 Evaluation
As noted by Okuyama et al. [28], no dataset exists for objective quantitative evaluation of body shape editing with large deformations. We therefore construct the first benchmark dataset for evaluating human body shape editing, consisting of 3,600 image pairs. The dataset uses distinct real face images and background descriptions to prevent training data leakage and includes diverse body shape variations for comprehensive assessment.
We evaluate generated images against ground truth using four key metrics. For image fidelity, we use Structural Similarity Index Measure (SSIM[41], Peak Signal-to-Noise Ratio (PSNR) for pixel-level accuracy, and Learned Perceptual Image Patch Similarity (LPIPS[46] for perceptual similarity. For body shape accuracy, we fit SMPL models to both ground truth and generated images using [37], then compute the Scale Corrected Per-Vertex Euclidean error in neutral (T-)pose (PVE-T-SC[38]. This metric quantifies how accurately the generated image preserves the target body shape.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRen et al.Okuyama et al.OursRefer to caption
Reference
Target 50kgTarget 170kgRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRen et al.Okuyama et al.OursRefer to caption
Reference
Target 50kgTarget 170kg
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRen et al.Okuyama et al.OursRefer to caption
Reference
Target 50kgTarget 170kg
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRen et al.Okuyama et al.OursRefer to caption
Reference
Target 50kgTarget 170kg
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRen et al.Okuyama et al.OursRefer to caption
Reference
Target 50kgTarget 170kg
Figure 7: Qualitative comparison of body shape editing for target weights of 50kg and 170kg. Compared to Ren et al. [33] and Okuyama et al. [28], our method produces more anatomically consistent body shapes and realistic clothing behavior, including proportionally adjusted hands, smoother fabric flow, better crease placement, and properly aligned shirt stripes.
In-the-wild Images. Fig. 6 demonstrates inference on in-the-wild images, where our method effectively handles diverse poses, backgrounds, and clothing while preserving person identity. In addition to SMPL target shapes, we provide textual prompts—“Make the person fatter,” “Make the person thinner,” or “Make the person muscular”—to explicitly guide the desired transformations.
Fig. 1 further demonstrates our model’s capability to perform diverse shape transformations. The model accurately follows SMPL depth maps to generate multiple variations of thinner and fatter versions from the reference image.
Qualitative Comparisons. Fig. 7 presents a qualitative comparison of our approach with previous state-of-the-art methods by Ren et al. [33] and Okuyama et al. [28]. The former controls transformation through warping strength while the latter uses input weight and height. For fair comparison, we select target beta parameters such that the final SMPL-derived weight matches that shown by Okuyama et al. [28]. Our results demonstrate more realistic transformations according to the target weight, as our model simultaneously adjusts overall body shape, limb proportions, and clothing, resulting in anatomically consistent and visually convincing modifications.
Quantitative Comparisons. Tab. 2 presents evaluation metrics on our benchmark dataset. Our model outperforms all other methods across all metrics, achieving a PVE-T-SC of 7.52 mm, indicating that our generated body shapes closely match ground-truth shapes.
We compare against Ren et al. [33], setting warping strength according to the weight difference between input and ground truth images for fair evaluation. Since the implementation of Okuyama et al. [28] is unavailable, quantitative comparison is limited to Ren et al.
We also evaluate FLUX.1 Kontext[dev] [2], a prompt-based editing model. We design specific prompts instructing the model to “Make the person fatter,” “Make the person thinner,” or “Make the person muscular” while specifying target weights. However, due to the lack of fine-grained editing controls, FLUX.1 Kontext[dev] achieves limited performance, as shown in Tab. 2.
Table 2: Comparison of our model Odo against Ren et al. [33] and FLUX.1 Kontext[dev] on our test set. We also show results for three ablations: without prompt conditioning in ReshapeNet, without ReferenceNet (IP-Adapter only), and training only on BR-5K data.
[…]
Ren et al. [33] 0.6790 17.4567 0.2363 13.6337
FLUX.1 0.6788 16.5195 0.2826 19.1911
Kontext[dev] [2]
Ours 0.7701 19.0080 0.2151 7.5214
Ours w/o 0.7281 15.1745 0.2566 9.8032
prompts
Ours w/o 0.6035 13.9982 0.4625 9.4157
ReferenceNet
Ours with 0.6939 14.1581 0.3432 18.6143
BR-5K data
5.3 Ablations
We conduct ablation experiments to evaluate the effectiveness of various modules in our model. Specifically, we investigate different methods for inputting image features into ReshapeNet and also explore different prompting techniques to guide the model in performing specific shape transformation tasks.
We evaluate our method without the ReferenceNet by removing it and relying solely on a trainable IP-Adapter to incorporate input image features into the denoising UNet. As shown in Fig. 8, omitting the ReferenceNet results in poor preservation of the background, attire (shoes in this example) and identity from the input image.
We also experiment with alternative prompts for the main denoising UNet. Specifically, instead of using category-specific prompts, we input a generic prompt such as “A photo of a person” for every pair during training. If only given this prompt during inference, the model remains effective at aligning with the SMPL depth maps and preserving clothing, background, and identity but struggles to enhance muscularity as shown in Fig. 8.
We also train our model only on the BR-5K dataset and evaluate it on our benchmark data to highlight the significance of our new dataset. Since the BR-5K dataset includes only very slight variations in body shape, the model is unable to learn to perform any significant body shape transformations as seen in Fig. 8. This is also reflected in the high value of the per-vertex error metric obtained on the test set as shown in Tab. 2.
The quantitative results for these ablations are shown in Tab. 2.
Figure 8: Omitting the category-specific prompts preserves identity, clothing, and background but fails to generate muscular transformations, despite aligning with the SMPL depth map. Removing the ReferenceNet allows the intended shape transformation but, loses identity and background details. With BR-5K data, our model is unable to perform any visible shape tranformation.
6 Conclusion
We introduce Odo, the first end-to-end diffusion-based framework for photorealistic, identity-preserving body shape editing at high resolution. Our approach generalizes across diverse images with varying poses and backgrounds by combining a ReferenceNet with a Depth ControlNet-conditioned diffusion model to produce fine-grained and diverse body shape modifications. We also introduce ChangeLing18K, a large-scale human body reshape dataset containing 18,573 paired images with varied transformations while maintaining consistent identity, clothing, and backgrounds.
Despite these advances, our model has certain limitations. It occasionally struggles to preserve fine facial details, leading to identity loss and facial artifacts. Additionally, inaccuracies during SMPL fitting—such as suboptimal parameter estimation from input images—can propagate errors to subsequent transformations. The model may also hallucinate body parts when target poses differ significantly from input poses. These limitations present opportunities for future work to enhance the fidelity and robustness of human body reshaping in natural images.